UT April 15
Installed motors for roof mirrors last week. telnet lbti-newubc 2225.
nic_mtr_mv <#> <steps>
2: right pupil, L-R. + mo...
AcquireAO is executed by WfsArbScripty.py and by the acquire module of source_acq.py. See those scripts for more details.
We've started an observing block for the LEECH survey this week. It's been a rocky beginning. We learned on Sunday of a coolant problem with the left...
Phils Notes for April 2013
[Download file "lbti_Feb2013_run_summary.pdf"]
Woke up yesterday afternoon to cloudy skies and some light snowfall. By midafternoon that had turned into windy conditions and heavy snowfall. The st...
January 31, 2013
Tracking LMIRCam and NOMIC temperature rise at Horizon.
LMIRCam goes up by 6 deg/hour from 60-70 K.
It was our first LBTI night of 2013, last night. The goal was to do phase sensing engineering the first half and exoplanet hunting (LEECH) the second...
Yesterday afternoon, a crew of four (Vanessa Bailey, Denis Defrere, Andy Skemer and myself) arrived on Mt. Graham to "wake up the instrument. Since l...
Scimeasure Firmware Compiler Files
[Download file "Compiler.zip"]
Request from HQ to Program Office:
[Download file "LBTI_PO_Request_2012-11-30.pdf"]
Charter for Implementation Review Board:
[Download file "LBTI_Oct-Nov2012RunSummary.pdf"]
Monday, Oct. 29
Putting Heaters in place at entrance to UBC
Took notes on position:
Left(SX): theta=285, x=-69 mm
NIC measurements - Nov. 2012
Phils Notes Oct-Nov. 2012
These are the measurements needed with NIC (not including on-sky tests) needed before turning the cooler off.
Powerpoint file with movies:
[Download file "Sep2012RunReport.zip"]
PDF file of above presentation:
[Download file "lbti_sep2012_run_sum...
SPC Stepper Motor
The SPC came up in a new configuration this run. It has a total range of approximately 21,000 steps. I think this means it...
We were on sky for our first of five nights in September commissioning,and carrying out early science with LBTI. The system was taken almost complete...